# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v6

@@ V-B2 Loco-manipulation tasks
While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it cannot physically interact with its environment. The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction. Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [2015_DAPRA_Challenge_1].
However, realising this opportunity is a significant challenge. Early studies, such as a ‘box transportation’ framework [2023_dao_OSUDRL_loco-manipulation_hybrid_simtoreal], often rely on inefficient, multi-policy solutions that lack visual perception. Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [2023_baltes_control_scooter, 2023_deepmind_haarnoja_playsoccer].
These difficulties create significant research opportunities. One such opportunity lies in exploring hierarchical control approaches. By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution. Using techniques like curriculum learning and imitation from human motion-capture data [2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining, 2023_seo_deep_loco_manipulation, 2024_Zhang_wholebody_adversarial_motion_priors, 2024_speratebody_expressive_allmotion], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
@@ V-B3 Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [quadModelBased1, quadModelBased2], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [quadDeformable]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [quadModular, humphreys2024bio], and sophisticated acrobatic motions are achieved through IL [quadImitation2].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [quadRLPlanning] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [quadRLLL, quadmpcref] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs an MPC to generate reference motions, which are then followed by an LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots module [TAMOLS] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are under way. These include mimicking real animals through motion-capture data and IL [quadAnimal1, quadAnimal2], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [quadManip1, quadManip2] or by using the robots’ legs [quadManip3]. Notably, the research presented in [quadManip2] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
@@ V-C1 Bipedal foundation models
Inspired by Robot Foundation Models (RFMs) [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture, 2023_generalpurposerobot_FMs], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion. A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments. Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots. Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [2023_VLA_Rt-2].
The proposed BFM paradigm would operate in two stages. First, IL on diverse datasets would establish the generalisable foundation. Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot. The potential of this approach is highlighted by recent works, with frameworks like FLaRe [2024_FLaRe_RLfinetune_foundationmodel] enhancing generalisation for long-horizon tasks, MOTO [2024_MOTO_RLfinetune_foundationmodel] enabling effective offline-to-online adaptation from images, and AdA [2023_RL_finetune_RMFs] demonstrating in-context adaptation to novel environments. Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
However, realising the BFM concept for bipeds presents significant challenges. The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [2024_FLaRe_RLfinetune_foundationmodel, 2023_RL_finetune_RMFs]. Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@ V-C2 Multi-layered adaptive models
A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [2022_DoAsICan, 2022_codeaspolicy_google] to parse commands into sub-tasks. For instance, Vision–Language Model Predictive Control [2024_VLM_PC] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
The mid-level planner selects or synthesises motions via learned motion libraries [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter] or DRL-based planners [2021_Mohammadreza_DRL_improveparameters_forplanner]. The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [2024_yangOu_LLM_hierarchical_quadrupedrobot], adaptive tracking controllers for whole-body tracking [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and imitation-based skills such as kicking and dancing [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task], by utilising RL, IL, or model-based methods. This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [2024_yangOu_LLM_hierarchical_quadrupedrobot].
However, realising the MLAM concept would introduce challenges distinct from BFMs. Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [TAMOLS]. Additionally, the computational latency incurred by HL reasoning modules like LLMs [2024_yangOu_LLM_hierarchical_quadrupedrobot] would pose limitations for tasks needing rapid reactions.
@@ VI Conclusion
Despite significant progress in DRL for robotics, a substantial gap remains between current achievements and the development of a unified framework capable of efficiently handling a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], climbing stairs [2021_siekmann_blind_DRL_stair], and navigating challenging terrains such as stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
@@ Appendix A Deep reinforcement learning algorithms
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-Layer Perceptrons (MLPs), a fundamental NN structure, excel in straightforward regression tasks with lower computational resource requirements. A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [2023_Rohan_sim2real_currentdecompose]. However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNNs) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diversity of NN architectures highlights the importance of selecting an appropriate model based on the specific requirements of bipedal locomotion tasks.
@@
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods, e.g. Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones], excel only in discrete action spaces and often struggle with high-dimensional action spaces. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to difficulties in accurate value estimation. In contrast, policy-based methods, such as policy-gradient techniques, can handle complex tasks but are generally less sample-efficient than value-based methods.
@@
More advanced algorithms combine both policy-based and value-based methods. The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [2015_continous_control_deep_reinforcement_learning, 2016_guided_learning_of_Control_Graphs_physics-based_characters]. Popular algorithms such as Trust Region Policy Optimisation (TRPO) [2015_TRPO_foundamentalpaper] and PPO, based on policy-based methods, borrow ideas from AC. Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [2023_DDPG_e2e_openaigym_mujoco], Twin Delayed Deep Deterministic Policy Gradients (TD3) [2019_stephen_TD3_instruction], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [2022_Jie_A3C_sim2real], and SAC (Soft Actor–Critic) [2022_Chen_transition_quadtobiped_TD3_SAC_Residual]. Each algorithm has its strengths for different tasks in bipedal locomotion scenarios. Several key factors determine their performance, such as sample efficiency, robustness and generalisation, and implementation complexity. A comparative analysis [2023_Omur_comparison_RLalgorithm_bipedallocomotion] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
@@
In [2017_Schulman_PPO_instructure], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC and DDPG [2023_Omur_comparison_RLalgorithm_bipedallocomotion]. Moreover, many studies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2020_Siekmann_drl_e2e_s2r_cassie] have demonstrated its robustness and ease of implementation. Combined with its flexibility to integrate with various NN architectures, this has made PPO the most popular choice in the field. Numerous studies have shown that PPO can enable the exploration of walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], jumping [2023_UCB_Cassie_RL_Jumping], stair climbing [2021_siekmann_blind_DRL_stair], and stepping-stone traversal [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark], demonstrating its efficiency, robustness, and generalisation.
@@
Additionally, the DDPG algorithm integrates the Actor–Critic framework with DQN to facilitate off-policy training, further optimising sample efficiency. In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2022_Chongben_pallel_DDPG_e23_silumation_only]. TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
@@
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences for training without requiring constant interaction with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [2022_Chen_transition_quadtobiped_TD3_SAC_Residual]. While A2C offers improved efficiency and stability compared with A3C, the asynchronous update mechanism of A3C provides better exploration capability and accelerates learning. Although these algorithms demonstrate clear advantages, they are more challenging to apply owing to their complexity compared with PPO.

### v7

@@ V-B2 Loco-manipulation tasks
While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it cannot physically interact with its environment. The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction. Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [4].
However, realising this opportunity is a significant challenge. Early studies, such as a ‘box transportation’ framework [22], often rely on inefficient, multi-policy solutions that lack visual perception. Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [7, 38].
These difficulties create significant research opportunities. One such opportunity lies in exploring hierarchical control approaches. By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [15].
Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution. Using techniques like curriculum learning and imitation from human motion-capture data [92, 112, 63, 97, 126, 19], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
@@ V-B3 Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [103, 114], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [20]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [27, 47], and sophisticated acrobatic motions are achieved through IL [32].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [33] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [51, 52] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs an MPC to generate reference motions, which are then followed by an LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots module [50] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are under way. These include mimicking real animals through motion-capture data and IL [85, 122], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [70, 30] or by using the robots’ legs [2]. Notably, the research presented in [30] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
@@ V-C1 Bipedal foundation models
Inspired by Robot Foundation Models (RFMs) [28, 44], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion. A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments. Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots. Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [128].
The proposed BFM paradigm would operate in two stages. First, IL on diverse datasets would establish the generalisable foundation. Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot. The potential of this approach is highlighted by recent works, with frameworks like FLaRe [43] enhancing generalisation for long-horizon tasks, MOTO [90] enabling effective offline-to-online adaptation from images, and AdA [8] demonstrating in-context adaptation to novel environments. Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
However, realising the BFM concept for bipeds presents significant challenges. The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [43, 8]. Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@ V-C2 Multi-layered adaptive models
A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [49, 65] to parse commands into sub-tasks. For instance, Vision–Language Model Predictive Control [18] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
The mid-level planner selects or synthesises motions via learned motion libraries [36, 61] or DRL-based planners [53]. The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [80], adaptive tracking controllers for whole-body tracking [31], and imitation-based skills such as kicking and dancing [83], by utilising RL, IL, or model-based methods. This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [80].
However, realising the MLAM concept would introduce challenges distinct from BFMs. Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [50]. Additionally, the computational latency incurred by HL reasoning modules like LLMs [80] would pose limitations for tasks needing rapid reactions.
@@ VI Conclusion
Despite significant progress in DRL for robotics, a substantial gap remains between current achievements and the development of a unified framework capable of efficiently handling a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [63], climbing stairs [99], and navigating challenging terrains such as stepping stones [25]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
@@ Appendix A Deep reinforcement learning algorithms
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-Layer Perceptrons (MLPs), a fundamental NN structure, excel in straightforward regression tasks with lower computational resource requirements. A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [102]. However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [98] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNNs) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diversity of NN architectures highlights the importance of selecting an appropriate model based on the specific requirements of bipedal locomotion tasks.
@@
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods, e.g. Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [74], excel only in discrete action spaces and often struggle with high-dimensional action spaces. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to difficulties in accurate value estimation. In contrast, policy-based methods, such as policy-gradient techniques, can handle complex tasks but are generally less sample-efficient than value-based methods.
@@
More advanced algorithms combine both policy-based and value-based methods. The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [66, 68]. Popular algorithms such as Trust Region Policy Optimisation (TRPO) [95] and PPO, based on policy-based methods, borrow ideas from AC. Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [45], Twin Delayed Deep Deterministic Policy Gradients (TD3) [21], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [57], and SAC (Soft Actor–Critic) [123]. Each algorithm has its strengths for different tasks in bipedal locomotion scenarios. Several key factors determine their performance, such as sample efficiency, robustness and generalisation, and implementation complexity. A comparative analysis [5] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
@@
In [96], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC and DDPG [5]. Moreover, many studies [98, 36, 100] have demonstrated its robustness and ease of implementation. Combined with its flexibility to integrate with various NN architectures, this has made PPO the most popular choice in the field. Numerous studies have shown that PPO can enable the exploration of walking [98], jumping [62], stair climbing [99], and stepping-stone traversal [25], demonstrating its efficiency, robustness, and generalisation.
@@
Additionally, the DDPG algorithm integrates the Actor–Critic framework with DQN to facilitate off-policy training, further optimising sample efficiency. In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [105, 106]. TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [123].
@@
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences for training without requiring constant interaction with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [123]. While A2C offers improved efficiency and stability compared with A3C, the asynchronous update mechanism of A3C provides better exploration capability and accelerates learning. Although these algorithms demonstrate clear advantages, they are more challenging to apply owing to their complexity compared with PPO.